We have the following date for the production of two producers, Juan and Maria. Juan can produce:
• 40 cups if he produces cups all day
• 40 plates if he produces plates all day Maria can produce:
• 10 cups if she produces cups all day
• 30 plates if he produces plates all day
Questions
1. DrawtheproductionpossibilityfrontiersofMariaandJuanusingMSword drawing tools you can find in the option Insert/Shapes (6 points)
2. Calculate the opportunity cost for cups and plates for Maria and Juan (6 points)
3. Calculate the Comparative advantage for cups and plates for Maria and Juan (6 points)
4. Whoshouldspecializeonwhat?(6points)
5. Findatradeagreementthatbothwouldacceptandshowthatbothwould
be better off (how many plates will exchange for 1 cup) (6 points)
1
Expert's answer
2019-11-08T06:03:17-0500


1.


2.When Juan producing 40 cups his opportunity cost is sacrificing 40 plates. Producing 40 plates opportunity cost is sacrificing 40 cups.When Maria producing 30 plates her opportunity cost is sacrificing 10 cups. producing 10 cups opportunity cost is sacrificing 30 plates.

3.

  • Juan: Comparative advantage of cups and plates"= (40-40)=0"
  • Maria: Comparative advantage of cups"= (10-30)=-20"
  •  Comparative advantage of plates"= (30-10)=20"


4.Since Juan producing 40 plates and 40 cups in all day he is the specialized in both productions.


5.If June can import 15 cups and 5 plates to Maria, then both can have total of 25 cups and 35 plates to make better trade agreement.